Use the five steps for problem solving to answer the following. Show all work.
Angle M and Angle N are supplementary angles. If angle M is nine more than one half Angle N, what are the measures of the angles?

Answer by edjones(8007)

m+n=180
m = n/2 + 9
2m=n+18 multiply each side by 2.
n=2m-18
m+2m-18=180 Substitute 2m-18 for n.
3m=198
m=66
n=114
.
Ed